# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ II.3 Control hardware
Minimizing quantum–classical feedback latency. The interaction between the quantum processor and classical control systems is characterized by quantum–classical feedback loops, which can be grouped into three primary timescales Ella et al. (2023). The first is the quantum real-time (QRT) feedback loop, executed within the coherence time of the qubits. An example of QRT feedback is active state preparation, where rapid response is critical to maintain qubit fidelity. Minimizing QRT feedback latency is essential because it directly influences the fidelity and overall performance of the quantum computer. The second feedback loop is the system real-time (SRT) loop, which occurs over one or more iterations beyond the qubit coherence time but is still dictated by the physical system. The SRT loop can be used to track qubit parameter drifts over time, and its speed impacts the system’s ability to adapt to noise and environmental fluctuations, thus affecting the quantum computer’s long-term fidelity. Lastly, the near real-time (NRT) feedback loop is used for classical-quantum interactions that are not constrained by immediate physical properties, but rather determined by the convergence time of quantum algorithms and calibration routines. NRT feedback often requires intensive classical computations and is typically carried out on the application level, outside the quantum control system. Efficient implementation of all these feedback loops is vital for both NISQ and QEC applications.
@@
Integration with high-performance computing resources. In addition to dedicated classical resources for tasks such as calibration, optimal control, and QEC decoding, the quantum workload may require high-performance compute resources. Examples for HPC use-cases include simulations and modeling, optimal control, and circuit knitting. Efficient integration with HPC requires a low-latency, high-throughput connectivity to HPC systems based on HPC standard interconnects such as Infiniband or Cray’s Slingshot or a dedicated protocol. Next, The control system should also be designed to process data locally, transmitting compact data streams to HPC servers and receiving compact instructions to maximize efficiency.
@@ III.1 Fault-tolerant circuit compilation
Fault-tolerant compilation of quantum algorithms is more complicated than that of classical computer programs because the final physical circuit depends on the specific noise characteristics of the quantum processor. It is commonly understood that the number of non-Clifford logical operations (e.g., the $T$ count of the algorithm) is a good indicator of the approximate cost of running the quantum algorithm. However, assembling the quantum program for exact physical circuits to run in hours, days, or even months on a quantum computer with millions of qubits and sophisticated coprocessors for control and decoding is much more involved.
An operating system for a fault-tolerant computer must therefore perform offline and real-time QPU and decoder characterization, modelling, and performance analysis, and incorporate this information into the compilation pipeline for FTQC execution. We consider three main modules for such a software suite, namely, the FTQC compiler, the emulator (including a noise profiler), and the assembler, details of which are discussed in the following section and summarized in Figure 1. An example of such a software suite, used to conduct our benchmarking studies in this work, is 1QBit’s TopQAD (Topological Quantum Architecture Design) toolkit 1QB Information Technologies (2024) (1QBit); Silva et al. (2024b, a).
At the highest level of abstraction, an FTQC compiler is responsible for circuit transpilation, decomposition, and parallelization of multi-qubit lattice surgeries on logical data qubits  1QB Information Technologies (2024) (1QBit); Silva et al. (2024b, a). At the lowest level, the emulator receives noise models from qubit arrays provided by various QCVV experiments to emulate fault-tolerant protocols at lower distances ($d<30$ typically) and extrapolates logical error rates to higher distances (sometimes 100 or more, depending on the algorithm; see Table 4). The results from the compiler and emulator are provided to the assembler, which is responsible for allocating various zones within the architecture’s layout (e.g., for magic state preparations at lower distances, distillation factories at increasing distances, and code growth and switching) and placement of logical qubits in the algorithm zone and scheduling lattice surgeries.
@@
Eventually, the procedures performed by the operating system, including compilation, emulation, and assembly, deliver the exact sequence of instructions for all the stabilizer measurement rounds, logical operator measurements, and conditional recovery operations to be performed by the 1–10M+ physical qubits system to the controller. This information is also provided to the decoder, since it must keep track of the logical protocols being executed (e.g., memory, teleportation, or code growth). We use this framework to conduct detailed resource estimations as presented in Section IV for real-world quantum chemistry problems as well. Furthermore, we study the sensitivity of the performance of the fault-tolerant quantum computer to various hardware parameters in Section III.3, which is helpful for guiding the design and fabrication of QPUs.
@@
Resource estimation analyses discussed in Section IV also provide profiles of all the independent lattice surgeries required to be performed on the concept architecture illustrated in Figure 14, and described in more detail in Refs. 1QB Information Technologies (2024) (1QBit); Silva et al. (2024b, a) and also in the appendix. Figure 15 shows an example histogram illustrating the sheer scale of independent decoding problems that must be solved by the decoders. The enormous problem sizes and decoding speed required for a successful execution of FTQC demands a tightly integrated high-performance decoding system. We describe such a decoding system in Section III.6.
Figure 15: Decoder requirements for electronic-structure quantum simulations of the $p$-benzyne molecule for an active space involving 26 6-31G basis set orbitals, using Trotterization based on the second-order Trotter–Suzuki product formula, with rigorous analytic error bounds (see Section IV). The histogram illustrates the scale of independent lattice surgery procedures that must be performed within the memory zone of the studied topologogical architecture to execute the quantum simulation circuits, with the top horizontal axis displaying the size of the independent decoding problems that must be solved by decoders. Independent decoding tasks require processing terabytes-large decoding graphs per second. Moreover, independent surgeries can involve 1M+ qubits spread across tens of DRs.
@@ III.2 Benchmarking quantum hardware for the quantum memory experiment
The results of our simulations are illustrated in Figure 16. We use numerical simulations at lower distances and extrapolate the logical infidelities at the higher distances in the regime of interest for utility-scale FTQC. To model an exponential logical error decay model, the extrapolation is based the model $\mu d^{2}\Lambda^{-(d+1)/2}$, where $d$ is the code distance and $\mu$ and $\Lambda$ are fitting parameters. We refer the reader to Ref. Silva et al. (2024a) for further details on the choice of this error suppression model. We note that previous works instead use the model $\mu\Lambda^{-(d+1)/2}$ to demonstrate an exponential suppression in the surface code error rates per cycle. Converting this per-cycle measure to an error model for the entire fault-tolerant protocol results in the model $\mu d\Lambda^{-(d+1)/2}$, which has a linear coefficient $d$ that is different from the coefficient we choose Fowler et al. (2012b); Acharya et al. (2023, 2024). To mitigate the bias introduced by small distances, data points with a logical infidelity below $10^{-2.5}$ are ignored in the fitting. The extracted $\Lambda$ value is an important hardware characteristic, as it determines the rate of logical error suppression with distance Chen et al. (2021). We note that the extracted $\Lambda$ value for baseline and target hardware parameters are, respectively, $2.34(1)$ and $9.3(3)$, showing an improvement by roughly a factor of $4$. The extracted value of $\Lambda$ for the $\Lambda_{18}$ model is $18(1)$, demonstrating an additional improvement factor of $2$ in the error suppression rate as compared to the target parameter set.
Figure 16: Hardware parameters benchmarked using a logical memory experiment. The numerical dataset is plotted alongside extrapolations derived from a two-parameter fit to $\mu d^{2}\Lambda^{-(d+1)/2}$, where $d$ is the code distance. The dataset is obtained using Clifford simulations based on the noise model described in this section. The extracted parameters, along with the fitting errors, are specified in the legend. The shaded region represents the regime of logical infidelities required to implement the electronic-structure simulations’ quantum circuits used in our QRE studies, presented in Section IV.
@@ III.3 Sensitivity of FTQC performance to specific hardware improvements
Quantum hardware engineers often face the uncertainty of which types of noise and errors have the most significant impact on the performance of FTQC, that is, of which hardware parameters are the most critical for achieving improved logical performance. It is unclear whether the coherence time of qubits or the two-qubit error rates matter most, or the state preparation and measurement (SPAM) errors are most crucial. In this section, we report results of a sensitivity analysis addressing this uncertainty.
We investigate the performance sensitivity of FTQC to specific hardware characteristics, as specified in Table 1, by assessing the logical error suppression factor $\Lambda$ as a function of individual hardware parameters. We analyze the following categories of hardware parameter improvements in the operations involved in implementing the quantum memory experiment: (i) coherence improvements (for idling physical qubits) involving $T_{1}$ and $T_{2}$ times; (ii) gate-control improvements affecting the Hadamard and CNOT gate infidelities; (iii) SPAM improvements concerning preparation, measurement, and reset errors; and (iv) a combined class encompassing all three groups. We determine the improvement in the error suppression rate $\Lambda$ when each of these parameter sets are improved separately, while keeping the others constant, as well as when all hardware parameters are improved simultaneously. Our findings are illustrated in  Figure 17. We observe that improvements in the gate-control errors yield the most significant impact, whereas improvements in SPAM errors and coherence enhancements are significantly less effective for achieving higher $\Lambda$ values. The $\Lambda$ value increase resulting from improving all parameters simultaneously is higher than the sum of individual increments of $\Lambda$. Our findings suggest that quantum gate fidelity improvements are more important than SPAM and idling qubit error rates for achieving greater logical performance.
@@ III.4 Impact of qubit and gate quality distributions on logical error rates
As our QRE studies in Section IV show, a utility-scale quantum computer is expected to require millions of qubits. Any manufacturing process that produces such large QPUs, or clusters of QPUs, will inevitably create qubits and gates of varying quality. In this section, we present analyses of the possible impacts of process variability on the logical error rates of the rotated surface code.
In order to obtain a realistic distribution of qubit and gate qualities, we use publicly available calibration data obtained for the ibm_torino quantum processor Quantum . In particular, we focus on the distributions of the $T_{1}$ times, as well as single-qubit, two-qubit, and readout errors. In Figure 18 we plot the cumulative distribution functions (CDF) of this data. Physically, we expect some correlations between these distributions, for example, longer coherence time for qubits should allow for higher fidelity or faster quantum control on the gates and therefore higher single-qubit and two-qubit fidelities.
@@
Figure 19: Correlations between the true and predicted single-qubit, two-qubit, and readout errors as a function of the $T_{1}$ time. The Pearson correlation coefficients $\rho$ are reported at the top of each figure.
@@
In Figure 20(a) we show the CDFs of three distributions generated by applying the transformation (1). The three values of $\sigma$ are chosen to represent different amounts of reductions in process variability from that of the studied QPU.
@@
    Simulate the fault-tolerant memory protocol using the assigned gate fidelities, and assuming gate times from the ibm_torino data (see Table 2), to determine a logical error rate for the code.

### v2

@@ II.3 Control hardware
Minimizing quantum–classical feedback latency. The interaction between the quantum processor and classical control systems is characterized by quantum–classical feedback loops, which can be grouped into three primary timescales Ella et al. (2023). The first is the quantum real-time (QRT) feedback loop, executed within the coherence time of the qubits. An example of QRT feedback is active state preparation, where rapid response is critical to maintain qubit fidelity. Minimizing QRT feedback latency is essential because it directly influences the fidelity and overall performance of the quantum computer. The second feedback loop is the system real-time (SRT) loop, which occurs over one or more iterations beyond the qubit coherence time but is still dictated by the physical system. The SRT loop can be used to track qubit parameter drifts over time, and its speed impacts the system’s ability to adapt to noise and environmental fluctuations, thus affecting the quantum computer’s long-term fidelity. Lastly, the near-real-time (NRT) feedback loop is used for classical–quantum interactions that are not constrained by immediate physical properties, but rather determined by the convergence time of quantum algorithms and calibration routines. Near real-time feedback often requires intensive classical computations and is typically carried out on the application level, outside the quantum control system. Efficient implementation of all these feedback loops is vital for both NISQ and QEC applications.
Control system integration with classical compute resources. By their nature, quantum computing workloads require quantum–classical iterations. These include calibration workflows, NISQ hybrid applications, and quantum error correction. Quantum–classical iterations require a low-latency, high-throughput interface with the right division of responsibilities that allows the transfer of a compact representation of the data across systems. For applications that require significant compute resources, tight integration to HPC clusters is required to efficiently utilize the quantum and classical resources. The integration between the control system and classical compute resources may be characterized by the feedback loop timescale. QRT feedback loops, executed within the coherence time of the qubits, require low-latency feedback and are therefore executed on the pulse processor, a dedicated real-time processor within the control system. For multi-qubit based feedback, the control system must support scalable, low-latency data distribution. SRT calculations may be performed by the control pulse processor or by classical resources connected to the control system such as CPUs or GPUs. SRT compute resources do not need to be physically located within the same rack as the quantum controller and may be connected through low-latency optical links. NRT calculations may require significant computational power, particularly for hybrid applications such as circuit knitting, and are typically performed on the client resources, outside the control system. The quantum control may be directly connected to a cluster high-performance network that provides scalable, high-bandwidth, and low-latency connectivity to the compute servers. In addition to hardware integration, we require a unified software framework that supports co-development, compilation, and execution of the quantum and classical portions of the application.
Efficient calibration. Calibration impacts qubit fidelity and plays a critical role in determining the error correction code distance, which in turn directly affects the scalability and complexity of the decoding process. A significant challenge is variability in qubit fidelity; the system’s overall performance is often constrained by the lowest-performing outlier qubits. To maximize the performance of these outliers, advanced calibration strategies such as optimal control, reinforcement learning, demonstrated in Figure 12, and model-based simulations may be used. To meet the fidelity goals for large quantum processors, it will be necessary to have scalable and efficient calibration routines that enable concurrent calibration across qubits. Reducing the calibration time allows to repeat the calibration more frequently and to allocate a larger portion of the time to optimizing outlier qubit performance. Figure 12 demonstrates the fidelity improvement from frequent calibration, tracking the drift in a qubit frequency. A key requirement for calibration flows is minimal execution overhead and feedback latency. A typical calibration node requires executing thousands or more shots. A shot is structured with initialization, executing a pulse sequence, and measurement, and typically takes hundreds of nanoseconds. As such, the calibration routine overhead should be minimized accordingly to few milliseconds from loading to gathering the measurements statistics.
Architecture of a quantum–classical integrated platform for FTQC. FTQC workflows require a quantum–classical integrated platform to execute the logic circuit, orchestrate QEC decoding and surface operations, and control physical qubits. The main components of the platform illustrated in Figure 13 include the quantum control, QEC acceleration servers, and dilution fridges that hold the quantum processor and cryogenic control. The quantum control is divided into clusters; typically each cluster controls a set of surfaces, and the independent clusters simplify the control and provide resiliency to errors while maintaining synchronized execution. The QEC acceleration servers are responsible for real-time decoding and logic circuit orchestration. The servers leverage general-purpose GPU/CPU systems with the option for additional ASIC or FPGA accelerators. The low latency QEC network provides an efficient interface for transferring syndromes and surface operations. A high-performance user network is used for data and program loading as well as connectivity to HPC resources required, for example, for logic circuit compilation.
Real-time calibration and control with AI on DGX Quantum. Characterizing, tuning, controlling, and optimizing quantum devices is a time-consuming process requiring various physical resources and particular expertise. AI can help automate these tasks, reducing experts’ time and effort while improving accuracy. Methods such as those based on neural networks and Bayesian optimization are particularly useful because they can produce accurate results from a small amount of data without relying on detailed physical models. Various AI methods have already been applied to characterize quantum devices Arshad et al. (2024); Craig et al. (2024); Berritta et al. (2024a); Wozniakowski et al. (2020), automate various tuning steps Schuff et al. (2024); Severin et al. (2024); Kalantre et al. (2019); Teske et al. (2019), optimize qubit control parameters such as gate fidelity and coherence times Sivak et al. (2022); Kurman et al. (2024a); Berritta et al. (2024b); Scerri et al. (2020); Arshad et al. (2024), and enhance qubit readout Seif et al. (2018); Phuttitarn et al. (2024). These methods can identify optimal settings more efficiently than traditional trial-and-error approaches. For example, neural networks can predict the ideal control parameters for qubits based on previous tuning data, while Bayesian optimization can be used to navigate the large parameter spaces involved in device calibration. The use of GPUs and AI infrastructure like those available with DGX Quantum (see Appendix F for details) is key to accelerating AI methods, and thus to real-time calibration and control.
Another important related concept is digital twins of quantum devices. A digital quantum twin can accurately capture the stochastic and non-Markovian behavior of various types of qubits. Using this approach, a digital quantum twin is built first, then used for characterizing, tuning, controlling, and optimizing real quantum devices. Building an AI version of a digital quantum twin would require a massive number of simulations for generating synthetic data for pre-training, performing pre-training, and inference. DGX Quantum NVIDIA Corporation (2025a), which combines GPUs, quantum integration through CUDA-Q NVIDIA Corporation (2025b), and the Nvidia AI framework, could accelerate these simulations significantly. The detailed description of DGX Quantum system and CUDA-Q platform are provided in Appendices F and G respectively.
@@
Integration with high-performance computing resources. In addition to dedicated classical resources for tasks such as calibration, optimal control, and QEC decoding, the quantum workload may require high-performance compute resources. Examples for HPC use-cases include simulations and modeling, optimal control, and circuit knitting. Efficient integration with HPC requires a low-latency, high-throughput connectivity to HPC systems based on HPC standard interconnects such as Infiniband or Cray’s Slingshot or a dedicated protocol. Next, The control system should also be designed to process data locally, transmitting compact data streams to HPC servers and receiving compact instructions to maximize efficiency. A key requirement for ensuring efficient HPC or cloud integration is co-scheduling of the resources to prevent delays caused by one resource waiting for others to become available.
@@ III.1 Fault-tolerant circuit compilation
Fault-tolerant compilation of quantum algorithms is more complicated than that of classical computer programs because the final physical circuit depends on the specific noise characteristics of the quantum processor. It is commonly understood that the number of non-Clifford logical operations (e.g., the $T$ count of the algorithm) is a good indicator of the approximate cost of running the quantum algorithm. However, assembling the quantum program for exact physical circuits to run in hours, days, or even months on a quantum computer with millions of qubits and sophisticated coprocessors for control and decoding is much more involved. A quantum OS for the fault-tolerant computer must therefore perform offline and real-time QPU and decoder characterization, modelling, and performance analysis, and incorporate this information into the compilation pipeline for FTQC execution. In what follows, we discuss three main modules for such an OS: the FTQC compiler, the emulator (including a noise profiler), and the assembler, as summarized in Figure 2.
At the highest level of abstraction, an FTQC compiler is responsible for circuit transpilation, decomposition, and parallelization of multi-qubit lattice surgeries on logical data qubits  1QB Information Technologies (2024) (1QBit); Silva et al. (2024b, a). At the lowest level, the emulator receives noise models from qubit arrays provided by various QCVV experiments to emulate fault-tolerant protocols at lower distances (typically $d<30$) and extrapolates logical error rates to higher distances (sometimes 100 or more, depending on the algorithm; see Table 5). The results from the compiler and emulator are provided to the assembler, which is responsible for allocating various zones within the architecture’s layout (e.g., for magic state preparations at lower distances, distillation factories at increasing distances, and code growth and switching) and placement of logical qubits in the algorithm zone and scheduling lattice surgeries.
@@
Eventually, the procedures performed by the OS, including compilation, emulation, and assembly, deliver the exact sequence of instructions for all the stabilizer measurement rounds, logical operator measurements, and conditional recovery operations to be performed by the 1–10M+ physical qubits system to the controller. This information is also provided to the decoder, since it must keep track of the logical protocols being executed (e.g., memory, teleportation, or code growth). We use this framework to conduct detailed resource estimations as presented in Section IV for real-world quantum chemistry problems as well. Furthermore, we study the sensitivity of the performance of the fault-tolerant quantum computer to various hardware parameters in Section III.3, which helps guiding the design and fabrication of QPUs.
@@
Resource estimation analyses discussed in Section IV also provide profiles of all the independent lattice surgeries required to be performed on the concept architecture illustrated in Figure 15, and described in more detail in Refs. 1QB Information Technologies (2024) (1QBit); Silva et al. (2024b, a) and also in Appendix B. Figure 16 shows an example histogram illustrating the sheer scale of independent decoding problems that must be solved by the decoders. The enormous problem sizes and decoding speed required for a successful execution of FTQC demands a tightly integrated high-performance decoding system. We describe such a decoding system in Section III.6.
Figure 16: Decoder requirements for electronic-structure quantum simulations of the $p$-benzyne molecule for an active space involving 26 6-31G basis set orbitals, using Trotterization based on the second-order Trotter–Suzuki product formula, with rigorous analytic error bounds (see Section IV). The histogram illustrates the scale of independent lattice surgery procedures that must be performed within the memory zone of the stopologogical architecture to execute the quantum simulation circuits, with the top horizontal axis displaying the size of the independent decoding problems that must be solved by decoders. Independent decoding tasks require processing terabytes-large decoding graphs per second. Moreover, independent surgeries can involve 1M+ qubits spread across tens of DRs.
@@ III.2 Benchmarking quantum hardware for the quantum memory experiment
The results of our simulations are illustrated in Figure 17. We use numerical simulations at lower distances and extrapolate the logical infidelities at the higher distances in the regime of interest for utility-scale FTQC. We regress the lowest-order term of the logical error model of the surface code
$\mu d^{2}\Lambda^{-(d+1)/2}$ (1)
to our numerical data, where $d$ is the code distance and $\mu$ and $\Lambda$ are fitting parameters. We refer the reader to Ref. Silva et al. (2024a) for further details on the choice of this error suppression model.
To mitigate the bias introduced by small distances, data points with a logical infidelity below $10^{-2.5}$ are ignored in the fitting. The extracted $\Lambda$ value is an important hardware characteristic, as it determines the rate of logical error suppression with distance Chen et al. (2021). We note that the extracted $\Lambda$ values for baseline and target hardware parameters are, respectively, $2.34(1)$ and $9.3(3)$, showing an improvement by roughly a factor of $4$. The extracted value of $\Lambda$ for the desired model is $18(1)$, demonstrating an additional improvement factor of $2$ in the error suppression rate as compared to the target parameter set.
We note that previous works Fowler et al. (2012b); Acharya et al. (2023, 2024) use the model $\bar{\mu}\bar{\Lambda}^{-(d+1)/2}$ to demonstrate an exponential suppression in the surface code error rates per cycle. Converting this per-cycle measure to an error model for the entire fault-tolerant protocol results in the model
$\bar{\mu}d\bar{\Lambda}^{-(d+1)/2},$ (2)
which has a linear coefficient $d$, as opposed to $d^{2}$ as in our model. In Figure 17, we show the discrepancies between the two choices in predicting logical error rates, especially at greater distances and with less-performant hardware specifications. The values of all fitting parameters for both models and our three hardware parameters are summarized in Table 2.
Figure 17: Prediction of high-distance error rates from numerical simulations at lower distances. The numerical dataset is plotted alongside extrapolations derived from the two-parameter fits to the models Equation 1 (shown using solid lines) and Equation 2 (shown using dashed lines). The dataset is obtained using Clifford simulations based on the noise model described in this section. The fitting parameters are provided in Table 2. The shaded region represents the regime of logical infidelities required to implement the electronic-structure simulations’ quantum circuits used in our QRE studies, presented in Section IV.
Baseline Target Desired
$\Lambda$ 2.34(1) 9.3(3) 18(1)
$\mu$ 0.0038(2) 0.019(4) 0.04(1)
$\bar{\Lambda}$ 2.119(6) 7.5(1) 13.5(3)
$\bar{\mu}$ 0.0259(8) 0.055(6) 0.082(9)
Table 2: Fitting parameters obtained for the models displayed in Equation 1 and Equation 2 when regressed to the numerical data of our three hardware parameter sets, specified as baseline, target, and desired.
@@ III.3 Sensitivity of FTQC performance to specific hardware improvements
It is often unclear a priori which types of noise and errors have the most significant impact on the performance of FTQC, and which hardware parameters are the most critical for achieving improved logical performance. It is unclear whether the coherence time of qubits or the two-qubit error rates matter most, or the state preparation and measurement (SPAM) errors are most crucial. In this section, we report results of a sensitivity analysis addressing this uncertainty.
We study the performance sensitivity of FTQC to specific hardware characteristics, as specified in Table 1, by assessing the logical error suppression factor $\Lambda$ as a function of individual hardware parameters. We analyze the following categories of hardware parameter improvements in the operations that implement the quantum memory experiment: (i) coherence improvements (for idling physical qubits) involving $T_{1}$ and $T_{2}$ times; (ii) gate-control improvements affecting the Hadamard and CNOT gate infidelities; (iii) SPAM improvements concerning preparation, measurement, and reset errors; and (iv) a combined class encompassing all three groups. We determine the improvement in the error suppression rate $\Lambda$ when each of these parameter sets are improved separately, while keeping the others constant, as well as when all hardware parameters are improved simultaneously. Figure 18 illustrates our findings. We observe that improvements in the gate-control errors yield the most significant impact, whereas improvements in SPAM errors and coherence enhancements are significantly less effective for achieving higher $\Lambda$ values. The $\Lambda$ value increase resulting from improving all parameters simultaneously is higher than the sum of individual increments of $\Lambda$. Our findings suggest that quantum gate fidelity improvements are more important than SPAM and idling qubit error rates for achieving greater logical performance.
@@ III.4 Impact of qubit and gate quality distributions on logical error rates
As our QRE studies in Section IV show, a utility-scale quantum computer is expected to require millions of qubits. Any manufacturing process that produces such large QPUs, or clusters of QPUs, will inevitably create qubits and gates of varying quality. In this section, we analyze possible impacts of process variability on the logical error rates of the rotated surface code.
In order to obtain a realistic distribution of qubit and gate qualities, we use publicly available calibration data obtained for the ibm_torino quantum processor Quantum . In particular, we focus on the distributions of the $T_{1}$ times, as well as single-qubit, two-qubit, and readout errors. In Figure 19 we plot the cumulative distribution functions (CDF) of this data. Physically, we expect some correlations between these distributions, for example, a longer coherence time for qubits should allow for higher fidelity or faster quantum control on the gates and therefore higher single-qubit and two-qubit fidelities.
@@
Figure 20: Correlations between the true and predicted single-qubit, two-qubit, and readout errors as a function of the $T_{1}$ time. The Pearson correlation coefficients $\rho$ are shown at the top of each figure.
@@
In Figure 21(a), we show the CDFs of three distributions generated by applying the transformation (3). The three values of $\sigma$ are chosen to represent different amounts of reductions in process variability from that of the studied QPU.
@@
    Simulate the fault-tolerant memory protocol with assigned gate fidelities, and assuming gate times from the ibm_torino device data (see Table 3), to determine the logical error rate for the code.